4. A mass of 2 kg is held 4 meters above the ground. The gravitational potential energy of this object, relative to the ground, is ___.

5. Which is easier, stopping a 1000 kg car traveling at 40 m / s or a 2000 kg car traveling at 20 m / s?

8. Suppose an airplane fires bullets in the same direction it travels. Does the momentum of the plane increase, decrease or remain the same? Explain
Answer #1

here,

4)

mass , m = 2 kg

height above the ground , h = 4 m

the gravitational potential energy of this object , PE = m * g * h

PE = 2 * 4 * 9.81 J = 78.48 J

5)

the impulse delivered by first car , I1 = 1000 * 40 kg.m/s = 4 * 10^4 kg.m/s

the impulse delivered by second car , I2 = 2000 * 20 kg.m/s = 4 * 10^4 kg.m/s

so, same effort is required for both cars

6)

as the bullet is fired in same direction

the momentum of plane is decreased